Write a net ionic equation for the overall reaction that occurs when aqueous solutions of hydrosulfuric acid (H2S) and sodium hydroxide are combined. Is this correct?

H2S + NaOH = H2O + Na2S unbalanced
H2S + 2 NaOH = 2 H2O + Na2S balanced

H2S(aq) + 2Na(aq) + 2 OH- (aq) = 2 Na (aq) + S 2- (aq) + 2H2O (l)
H2S(aq) + 2 OH- (aq) = S 2- (aq) + 2H2O (l)
1
Expert's answer
2015-10-06T08:17:30-0400
H2S + NaOH = H2O + Na2S unbalanced
H2S + 2 NaOH = 2 H2O + Na2S balanced

H2S(aq) + 2Na(aq) + 2 OH- (aq) = 2 Na (aq) + S2- (aq) + 2H2O (l)
H2S(aq) + 2 OH- (aq) = S2- (aq) + 2H2O (l)
Answer:
Yes, these equations are correct.
Sodium hydroxide and sodium sulfide are electrolytes, so they form ions in water solution. Hydrosulfuric acid (H2S) isn’t electrolyte, so it doesn’t form ions in water solution.